OpenStudy (anonymous):

solution for x2-2x=6?

OpenStudy (anonymous):

\[x^2-2x=6\] is this it?

OpenStudy (anonymous):

yes! i need the solution or solutions

OpenStudy (anonymous):

ok so you are gonna bring over your 6 \[x^2 -2x-6=0\] looking at this i think we have to use the quadratic formula. Do you know how?

OpenStudy (anonymous):

not really good, i'm working on it

OpenStudy (anonymous):

Thats alright I am here to help :) so remember this \[ax^2 +bx+c\] \[\frac{ -b \sqrt{b^2 -4ac} }{ 2a }\] plugging in what we have makes it \[\frac{ -(-2)\sqrt{(-2)^2-4(1)(-6)} }{ 2(1)}\]

OpenStudy (anonymous):

opps i made a mistake there should be a + and - between (-b) and the square root

OpenStudy (anonymous):

so \[\frac{ 2\pm \sqrt{28} }{ 2 }\] so breaking it up \[\frac{ 2+2\sqrt{7} }{ 2}\] \[\frac{ 2-2\sqrt{7} }{ 2}\] because there is a 2 in front and a 2 in front of the radical it can be divided by the 2 on the bottom making them \[1+\sqrt{7}\] and \[1-\sqrt{7}\] Does this make sense?